OpenStudy (magbak):

I will award medal !!!!!!!!!!!!!!!!!!!!! Please help!!!!!!!!!!!!!!!!!!!! Using complete sentences, describe how you would check the quotient of two rational expressions. Create and demonstrate your own example of this concept.

OpenStudy (anonymous):

To check the quotient of a division problem, I would multiply the quotient by the divisor and add the remainder. If the quotient is correct, the result will be the dividend. \[\frac{ x ^{2}-x-6 }{ x-3 }\] = \[\frac{ (x-3) (x+2) }{ x-3 }\]=x+2 To check this, I will multiply the quotient which is x+2 by the divisor which is x-3. It the result is correct, I with get \[x ^{2}-x-6\] \[(x+2)(x-3)=x ^{2}-3x+2x-6=x ^{2}-x-6\]